OpenStudy (anonymous):

How many solutions does this system have? -3x+6y=10 and -3x+6y=-4 a. 0 b. 1 c. infinite d. 2 Need help it is due today! I will give medal for help!

OpenStudy (danjs):

welcome to open study!

OpenStudy (danjs):

-3x + 6y = 10 -3x + 6y = -4

OpenStudy (anonymous):

thanks:)

OpenStudy (danjs):

Before doing this problem, here are the possible choices. Zero solutions - The lines do not intersect. They must be parallel.

OpenStudy (anonymous):

yeah

OpenStudy (danjs):

1 solution - the lines intersect at 1 point (x,y) infinite solutions - both equations describe the same line, they overlap each other at infinite points

OpenStudy (danjs):

We can throw away choice D) 2 lines can never intersect at 2 points only

OpenStudy (anonymous):

yep i remember those things. i just need to know what it was because my teacher did not describe it well

OpenStudy (danjs):

ok

OpenStudy (danjs):

-3x + 6y = 10 -3x + 6y = -4 A little trick is, if the left hand sides are the same, but the right hand sides are different, like here, the lines are parallel

OpenStudy (anonymous):

well the -3x are the same, and the 6y are the same. But you said the right side has to be different. So do you mean the 10 and -4 or the 6y?

OpenStudy (danjs):

or, if you can multiply one equation by a constant number to make the left side the same , it is still parallel, like here 2x + 4y = ... 4x + 8y = .... multiply the first equation by 2, and you get the second , the two are parallel

OpenStudy (anonymous):

so they are parallel? ok. so what would the answer be?

OpenStudy (danjs):

The solution to a system of 2 lines is where they intersect, how many solutions do 2 parallel lines have?

OpenStudy (anonymous):

im not sure exactly but i think it would have two or 0

OpenStudy (danjs):

Parallel lines never intersect. They will have zero solutions / intersection points.

OpenStudy (anonymous):

ok. thank you! i will tell you if it is right!:)

OpenStudy (danjs):

Two lines can never have 2 solutions, they can't intersect 2 times. Try to draw two lines on paper that only intersect 2 times

OpenStudy (anonymous):

can you possibly help me with 3 more?

OpenStudy (danjs):

yes, post a new thread, type @DanJS in the question

OpenStudy (anonymous):

ok